# Comparative study of the butterfly velocity in holographic QCD models at finite temperature and chemical potential

### v2

@@ 4.2 $v_{B}$ from the entanglement wedge method
To compute $v_{B}$ using the entanglement wedge method, we first need to recast Eq. (4.1) in terms of Eq. (2.1). A straightforward comparison gives:
@@
In the above equation, $v_{B}^{2}$ has been decomposed into two parts: one that explicitly depends on the chemical potential and the other, which is independent of it. The separated expressions clearly indicate that a nontrivial dependence of $v_{B}^{2}$ on $T$ and $\mu$ is expected in this model. Since it is not possible to analytically invert (4.1) to express $z_{h}$ in terms of $\mu$ and $T$, we employ numerical methods to study $v_{B}^{2}$ as a function of these parameters. As mentioned above, this holographic model enjoys the Hawking-Page phase transition, with the black hole phase being thermodynamically favored only at high temperature. Accordingly, below we mainly concentrate on the thermodynamically stable black hole phase with temperature $T\geq T_{c}$, corresponding to a deconfined phase in the dual boundary theory.
@@
Our results for the butterfly velocity as a function of chemical potential and temperature are shown in Figure 5. Here we concentrate on the temperature range $T=T_{c}$ to $T=3.0~{}T_{c}$ and chemical potential range $\mu=0$ to $\mu=1.0$; which are relevant for the deconfined phase of QCD. We find that $v_{B}^{2}$ decreases monotonically with the chemical potential in the deconfined phase. This is true for all temperatures greater than the deconfined temperature. Similarly, for a fixed $\mu$, $v_{B}^{2}$ increases monotonically with temperature at low temperatures and then saturates to a constant value at high temperatures. Again, this constant value is $2/3$ and is independent of the chemical potential. This result can also be analytically observed from Eq. (4.9). Note that in the large temperature limit $T\rightarrow\infty$, corresponding to $z_{h}\rightarrow\infty$, we get
$v_{B}^{2}=\frac{2}{3}-\frac{6a+\mu^{2}}{9z_{h}^{2}}+\mathcal{O}(z_{h}^{-4})\,,$ (4.10)
suggesting that model parameters and chemical potential only give subleading corrections to $v_{B}^{2}$ in the large temperature limit. The above behavior of $v_{B}^{2}$ is quite similar to its behavior in the 1RCBH and 2RCBH models discussed earlier, i.e., the rate at which information about the perturbation propagates among the local degrees of freedom in the dual boundary deconfined phase decreases/increases with the chemical potential/temperature.
@@ 4.4 $v_{B}$ from the pole-skipping method
Matter Lagrangian of the holographic model considered in this section is the same as (3.27) of the 1RCBH and 2RCBH models considered in the previous section, as both models are constructed from the same EMD action. This implies that the Eq. (2.44) also holds true for this model. Accordingly, the pole-skipping points are again given by Eq. (2.47). Substituting the metric functions from (LABEL:eqn:metricfunctionsourmodel) into (2.47), we obtain the following expression of the pole skipping points in the current holographic model:
@@
(b) Im($\omega_{*}$) as a function of $\mu$ and $T/T_{c}$
Figure 6: Pole-skipping points $q_{*}$ and $\omega_{*}$ as a function of temperature and chemical potential
@@ 5 Numerical Bottom-up Holographic QCD model
5.0.1 Background
In this section, we turn our attention to another bottom-up holographic QCD model, particularly to the model of [55], to probe the temperature and chemical potential dependent profile of the butterfly velocity. Again, this model is based on the EMD gravity (3.1). Unlike the EMD models considered in the previous sections, this model is completely numerical and requires numerical integration of the gravity equations of motion from the horizon to the boundary to obtain boundary field theory information. Since the foundational work of [54], numerous numerical EMD-based AdS/QCD models have been developed, including [55], and others such as [94, 95, 96, 97, 98, 99]. These holographic models share a common strategy: the dilaton potential $V(\phi)$ and the gauge kinetic function $\tilde{f}(\phi)$ are chosen to match lattice QCD results for observables like entropy density, pressure, and susceptibility at zero chemical potential. Once the forms of $V(\phi)$ and $\tilde{f}(\phi)$ are calibrated at zero chemical potential, results with nonzero chemical potential constitute genuine holographic predictions. Over the past decade, such numerical models have become key tools for exploring QCD via holography. In what follows, we primarily examine the model of [55], though our analysis of the butterfly velocity can be readily extended to other similar frameworks.
Before we explicitly analyze the butterfly velocity in the model of [55], let us first briefly discuss its underlying mathematical structure and numerical implementation. In this bottom-up EMD model, the following form of the dilaton potential and coupling function is considered
@@ 5.0.1 Background
with $c_{1}=0.606$, $c_{2}=0.703$, $c_{4}=-0.1$, $c_{6}=0.0034$, $b_{1}=1.2$, $b_{2}=0.69$, and $b_{3}=100$. The free parameters $c_{i}$ and $b_{i}$ are again chosen to match with physical properties of realistic QCD. The Ansätze
$ds^{2}=e^{2\eta(z)}\left[-\mathcal{G}(z)dt^{2}+dx_{1}^{2}+dx_{2}^{2}+dx_{3}^{2}\right]+\frac{e^{2\tau(z)}dz^{2}}{\mathcal{G}(z)}\,,$
$\phi=\phi(z)\,,~{}~{}~{}\mathcal{A}_{M}dx^{M}=\Phi(z)dt\,,$ (5.2)
represents an asymptotically $AdS_{5}$ spacetime with boundary at $z\rightarrow\infty$ and defines a black hole horizon by $\mathcal{G}(z_{h})=0$. Following [54], here we mainly work in the gauge $\tau(z)=0$. With Ansätze (5.0.1) for the metric, dilaton, and gauge fields, one gets the following four independent equations of motion
@@
The coordinates used in Eq. (5.0.1) are useful to numerically solve the EMD equations of motion. To make a meaningful comparison of solutions and calculate physical observables of dual theory, one should ideally arrange for the source terms $\phi_{A}$ to take the same value. This is generally done through a coordinate transformation [54]. In particular, suppose we numerically obtain a solution characterized by a positive value of $\phi_{A}$. We then aim to apply a coordinate transformation to this solution to express it in the following form
@@
$\tilde{\eta}(\tilde{z})=\tilde{z}+\mathcal{O}(e^{-2\nu\tilde{z}})\,,$
$\tilde{\mathcal{G}}(\tilde{z})=1+\tilde{\mathcal{G}}_{4}^{\text{far}}e^{-4\tilde{z}}+\mathcal{O}(e^{-(4+2\nu)\tilde{z}})\,,$
$\tilde{\Phi}(\tilde{z})=\tilde{\Phi}_{o}^{\text{far}}+\tilde{\Phi}_{2}^{\text{far}}e^{-2\tilde{z}}+\mathcal{O}(e^{-(2+\nu)\tilde{z}})\,,$
$\tilde{\phi}(\tilde{z})=e^{-\nu\tilde{z}}+\mathcal{O}(e^{-2\nu\tilde{z}})\,.$ (5.12)
To distinguish these coordinates, following [55], we refer to the coordinates with a tilde as standard coordinates, and those without a tilde as numerical coordinates. Note that the blackening function $\tilde{\mathcal{G}}(\tilde{z})$ goes to one at the asymptotic boundary in the standard coordinates, thereby allowing us to compute physical quantities, such as the temperature or chemical potential, using the standard holographic techniques. Now, setting $ds^{2}=d\tilde{s}^{2}$, $\tilde{\phi}(\tilde{z})=\phi(z)$, $\tilde{\Phi}(\tilde{z})d\tilde{t}=\Phi(z)dt$, and comparing the near-boundary expansion [Eq. (5.10) and (5.12)], one gets the following relations between the standard and numerical coordinates
@@
$\tilde{\eta}(\tilde{z})=\eta(z)-\log(\phi_{A}^{1/\nu})\,,~{}~{}~{}~{}\tilde{\mathcal{G}}(\tilde{z})=\frac{1}{\mathcal{G}_{0}^{\text{far}}}\mathcal{G}(z)\,,~{}~{}~{}~{}\tilde{\Phi}(\tilde{z})=\frac{1}{\phi_{A}^{1/\nu}\sqrt{\mathcal{G}_{0}^{\text{far}}}}\,\Phi(z),$ (5.14)
@@
Using the above relations, one can recast the thermodynamic quantities of the dual boundary theory, such as the temperature $T$, chemical potential $\mu$, entropy density $S$, charge density $\rho$, etc, in terms of the coefficients of the near-boundary expansions in the numerical coordinates
@@ 5.1 $v_{B}$ from the entanglement wedge method
The numerical results of the butterfly velocity as a function of temperature and chemical potential are shown in Figure 7. To obtain this result, we numerically constructed approximately $\sim 5\times 10^{5}$ charged black hole solutions across a two-dimensional grid of initial conditions $\{\phi_{0},\Phi_{1}\}$. The parameter $\phi_{0}$ was varied in uniform increments from $0.55$ to $5.5$, while for each value of $\phi_{0}$, the ratio $\Phi_{1}/\Phi_{1}^{\text{max}}(\phi_{0})$ was varied from 00 to $0.5$ in equally spaced steps. [^4] For this system, the maximum value $\Phi_{1}$ can take is bounded by $\phi_{0}$ and is given by $\Phi_{1}<\Phi_{1}^{max}\sqrt{-\frac{2V(\phi_{0})}{f(\phi_{0})}}$. We again remind that $\phi_{0}$ corresponds to the value of the scalar field at the horizon and $\Phi_{1}$ is the electric field at the horizon. We find that for relatively large temperatures, $T>200~{}\text{MeV}$, $v_{B}$ monotonically increases with the temperature, and then, like in the case of the other three models, it asymptotes to the conformal chargeless plasma value in the infinite temperature limit. However, as opposed to the other three models, $v_{B}$ now also shows a non-monotonic variation at relatively low temperature $T<130~{}\text{MeV}$, i.e., $v_{B}$ decreases with temperature for $T<130~{}\text{MeV}$.
At this point, it is important to keep in mind that, unlike the model considered in section 4, there is no Hawking/Page phase transition in the gravity side in the current holographic model. This hinders the determination of the dual deconfinement transition temperature via direct methods in this model. To get the deconfinement temperature, one therefore relies on indirect methods, such as the temperature at which the speed of sound attains a minimum value. From this indirect method, the deconfinement transition temperature is found to be around $T_{c}\sim 140~{}\text{MeV}$ [55], i.e., the transition temperature occurs beyond the point of nonmonotonicity. Therefore, it is safe to say that the butterfly velocity again exhibits monotonic thermal structure in the high temperature deconfined phase.
Furthermore, interestingly, just like in the case of the other three models, $v_{B}$ monotonically decreases with the chemical potential for all fixed values of the temperature. This observation further suggests a universal behavior in the rate of information propagation, associated with the spread of perturbations and operator growth, within the local degrees of freedom of the dual deconfined plasma, which decreases with increasing chemical potential. Note that due to the marginal decrease in $v_{B}$ with respect to $\mu$, this fact is not clearly evident from Figure 7, but one can verify this by obtaining a two-dimensional plot of $v_{B}$ vs $\mu$ at various fixed values of $T$.
@@
However, the behavior at low temperatures, particularly near the deconfinement temperature, is less transparent. Unlike in the UV, there is no general symmetry argument that dictates the observed similarity between the butterfly velocity and the speed of sound in the infrared (IR). We also analyze other numerical holographic QCD models of the type [55], such as [54] and [95], and find analogous thermal behavior between $v_{B}$ and $c_{s}$. This apparent parallel thermal behavior may suggest deeper dynamical connections or shared sensitivity to the underlying non-perturbative physics of the strongly coupled plasma, though further investigation is required to elucidate the mechanisms responsible for this parallel behavior.
@@ 5.2 $v_{B}$ from the OTOC method
$A(\tilde{u},\tilde{v})$ $=$ $\frac{\tilde{\mathcal{G}}(\tilde{z}_{h}-\tilde{u}\tilde{v})e^{2\tilde{\eta}(\tilde{z}_{h}-\tilde{u}\tilde{v})}}{2\alpha^{2}\tilde{u}\tilde{v}}\,,$
$B(\tilde{u},\tilde{v})$ $=$ $e^{2\tilde{\eta}(\tilde{z}_{h}-\tilde{u}\tilde{v})}\,.$ (5.19)
Computing the near-horiozn limits of $A$ and $\partial_{\tilde{u}}\partial_{\tilde{v}}B$ as $\tilde{u}\rightarrow 0$, we arrive at the following
@@ 5.3 $v_{B}$ from the pole-skipping method
Although analytic results for the energy-momentum tensor are difficult to obtain in the holographic QCD model considered in this section, it is straightforward to see that, as this model is diffeomorphism invariant and falls in the EMD gravity definition of type (3.1), the energy-momentum tensor again satisfies the condition (2.44). Accordingly, the pole-skipping points are again given by Eq. (2.47). Explicitly, after a little bit of algebra, they take the following form
@@
$v_{B}^{2}$ $=$ $\frac{\omega_{*}^{2}}{q_{*}^{2}}=\frac{1}{6\eta^{\prime}(z_{h})\mathcal{G}_{0}^{\text{far}}}\,,$ (5.24)
@@
We conclude by noting a technical subtlety in the holographic model considered in [55]. The last term in the gauge kinetic function $\tilde{f}(\phi)$, as defined in Eq. (5.1), becomes effectively negligible for most values of $\phi$ due to the large numerical value of the parameter $b_{3}$, contributing meaningfully only in the limit of very small $\phi$. This exponential term is included to ensure that $\tilde{f}(0)=1$. However, this choice also influences the scaling of the chemical potential and charge, particularly when fitting to susceptibility data (from which the parameters of $\tilde{f}(\phi)$ are fixed), and should therefore be handled with care. Enforcing $\tilde{f}(0)=1$ results in a spurious narrow spike in the function $\tilde{f}(\phi)$ at small $\phi$. This issue has been discussed in great detail in [99], where it was further pointed out that the physical impact of this spike is essentially negligible.

### v3

@@ 4.2 $v_{B}$ from the entanglement wedge method
To compute $v_{B}$ using the entanglement wedge method, we first need to recast Eq. (4.1) in terms of Eq. (2.1). A straightforward comparison gives
@@
In the above equation, $v_{B}^{2}$ has been decomposed into two parts: one that explicitly depends on the chemical potential and the other, which is independent of it. The separated expressions clearly indicate that a nontrivial dependence of $v_{B}^{2}$ on $T$ and $\mu$ is expected in this model. Since it is not possible to analytically invert (4.1) to express $z_{h}$ in terms of $\mu$ and $T$, we employ numerical methods to study $v_{B}^{2}$ as a function of these parameters. As mentioned above, this holographic model enjoys the Hawking/Page phase transition, with the large black hole phase being thermodynamically favored only at high temperature. Accordingly, below we mainly concentrate on the thermodynamically stable large black hole phase with temperature $T\geq T_{c}$, corresponding to a deconfined phase in the dual boundary theory.
Our results for the butterfly velocity as a function of chemical potential and temperature are shown in Figure 6. Here we concentrate on the temperature range $T=T_{c}$ to $T=3.0~T_{c}$ and chemical potential range $\mu=0$ to $\mu=1.0$, which are relevant for the deconfined phase of QCD. We find that $v_{B}^{2}$ decreases monotonically with the chemical potential in the deconfined phase. This is true for all temperatures greater than the deconfined temperature. Similarly, for a fixed $\mu$, $v_{B}^{2}$ increases monotonically with temperature at low temperatures and then saturates to a constant value at high temperatures. Again, this constant value is $2/3$ and is independent of the chemical potential. This result can also be analytically observed from Eq. (4.9). Note that in the large temperature limit $T\rightarrow\infty$, corresponding to $z_{h}\rightarrow\infty$, we get
$v_{B}^{2}=\frac{2}{3}-\frac{6a+\mu^{2}}{9z_{h}^{2}}+\mathcal{O}(z_{h}^{-4})\,$, (4.10)
suggesting that model parameters and chemical potential only give subleading corrections to $v_{B}^{2}$ in the large temperature limit. The above behavior of $v_{B}^{2}$ is quite similar to its behavior in the 1RCBH and 2RCBH models discussed earlier; i.e., the rate at which information about the perturbation propagates among the local degrees of freedom in the dual boundary deconfined phase decreases/increases with the chemical potential/temperature. Moreover, we have analyzed the butterfly velocity over a broad region of the temperature-chemical potential parameter space and consistently found that its behavior aligns with the pattern described above.
@@ 4.4 $v_{B}$ from the pole-skipping method
The matter Lagrangian of the holographic model considered in this section is the same as (3.27) of the 1RCBH and 2RCBH models considered in the previous section, as both models are constructed from the same EMD action. This implies that the Eq. (2.44) also holds true for this model. Accordingly, the pole-skipping points are again given by Eq. (2.47). Substituting the metric functions from (LABEL:eqn:metricfunctionsourmodel) into (2.47), we obtain the following expression of the pole-skipping points in the current holographic model:
@@
(b) Im($\omega_{*}$) as a function of $\mu$ and $T/T_{c}$.
Figure 7: Pole-skipping points $q_{*}$ and $\omega_{*}$ as a function of temperature and chemical potential.
@@ 5 Numerical Bottom-up Holographic QCD model
5.1 Background
In this section, we turn our attention to another bottom-up holographic QCD model, particularly to the model of [56], to probe the temperature- and chemical-potential-dependent profile of the butterfly velocity. Again, this model is based on the EMD gravity (3.1). Unlike the EMD models considered in the previous sections, this model is completely numerical and requires numerical integration of the gravity equations of motion from the horizon to the boundary to obtain boundary field theory information. Since the foundational work of [55], numerous numerical EMD-based AdS/QCD models have been developed, including [56], and others, such as [51, 106, 90, 107, 108, 109]. These holographic models share a common strategy: the dilaton potential $V(\phi)$ and the gauge-kinetic function $\tilde{f}(\phi)$ are chosen to match lattice QCD results for observables like entropy density, pressure, and susceptibility at zero chemical potential. Once the forms of $V(\phi)$ and $\tilde{f}(\phi)$ are calibrated at zero chemical potential, results with nonzero chemical potential constitute genuine holographic predictions. Over the past decade, such numerical models have become key tools for exploring QCD via holography. In what follows, we primarily examine the model of [56], though our analysis of the butterfly velocity can be readily extended to other similar frameworks.
Before we explicitly analyze the butterfly velocity in the model of [56] and other variants of it, let us briefly differentiate it from the models considered previously. Apart from the obvious difference at the solution level, which is now completely numerical as opposed to the previous analytical models, this model is mainly suitable for studying finite-temperature QCD features in the deconfined phase. In particular, the gravity equations are solved to construct numerical black holes in the bulk. However, setting up the numerical procedure to compute confined properties requires implementing the initial value problem deep in the bulk (strictly speaking, at $z=0$), which poses serious numerical challenges. Consequently, one faces problems in studying important observables, such as the Wilson and Polyakov loops, in the confined phase here. [^4] It is worth noting that, by choosing an appropriate dilaton potential, one can construct a four-dimensional gauge theory with a nontrivial renormalization group (RG) flow at zero temperature, which runs from an ultraviolet (UV) fixed point to an infrared (IR) fixed point. In the dual gravity description, this RG flow corresponds to a domain-wall geometry that interpolates between two AdS spaces [110]. This is in contrast to the previous models, where these observables can be computed explicitly. Similarly, unlike the holographic model of Section 4, it is difficult to directly obtain the deconfinement temperature here, as the gravity model does not exhibit the Hawking/Page phase transition. Accordingly, one can not consistently define the deconfinement temperature and rely on indirect techniques for the same.
Despite these limitations, such numerical models have become central to phenomenological studies of QCD via holography. They have been employed to determine the QCD critical point [106], study anisotropic heavy-quark drag forces and Langevin diffusion coefficients [90], analyze the entanglement and complexity structure of QCD phases [107, 111], and explore the behavior of QCD in the presence of a magnetic field [112], among other applications. More recently, these models have also been used to construct the QCD equation of state via neural networks [113], and to investigate stochastic gravitational waves and primordial black hole (PBH) production in the early Universe [114]. Here, we use such models to analyze the chaotic aspect of QCD.
Now we briefly discuss its underlying mathematical structure and numerical implementation. In this bottom-up EMD model, the following form of the dilaton potential and coupling function is considered:
@@ 5.1 Background
with $c_{1}=0.606$, $c_{2}=0.703$, $c_{4}=-0.1$, $c_{6}=0.0034$, $b_{1}=1.2$, $b_{2}=0.69$, and $b_{3}=100$. The free parameters $c_{i}$ and $b_{i}$ are again chosen to match with physical properties of realistic QCD. The ansatz
$ds^{2}=e^{2\eta(z)}\left[-\mathcal{G}(z)dt^{2}+dx_{1}^{2}+dx_{2}^{2}+dx_{3}^{2}\right]+\frac{e^{2\tau(z)}dz^{2}}{\mathcal{G}(z)}\,$, (5.2)
$\phi=\phi(z)\,,~~~\mathcal{A}_{M}dx^{M}=\Phi(z)dt\,$,
represents an asymptotically $AdS_{5}$ spacetime with boundary at $z\rightarrow\infty$ and defines a black hole horizon by $\mathcal{G}(z_{h})=0$. Following [55], here we mainly work in the gauge $\tau(z)=0$. With ansatz (5.2) for the metric, dilaton, and gauge fields, one gets the following four independent equations of motion:
@@
The coordinates used in Eq. (5.2) are useful to numerically solve the EMD equations of motion. To make a meaningful comparison of solutions and calculate physical observables of dual theory, one should ideally arrange for the source terms $\phi_{A}$ to take the same value. This is generally done through a coordinate transformation [55]. In particular, suppose we numerically obtain a solution characterized by a positive value of $\phi_{A}$. We then aim to apply a coordinate transformation to this solution to express it in the following form:
@@
[…]
@@
$\tilde{\eta}(\tilde{z})=\eta(z)-\log(\phi_{A}^{1/\nu})\,,~~~~\tilde{\mathcal{G}}(\tilde{z})=\frac{1}{\mathcal{G}_{0}^{\text{far}}}\mathcal{G}(z)\,,~~~~\tilde{\Phi}(\tilde{z})=\frac{1}{\phi_{A}^{1/\nu}\sqrt{\mathcal{G}_{0}^{\text{far}}}}\,\Phi(z)\,.$ (5.14)
@@
Using the above relations, one can recast the thermodynamic quantities of the dual boundary theory, such as the temperature $T$, chemical potential $\mu$, entropy density $S$, charge density $\rho$, etc., in terms of the coefficients of the near-boundary expansions in the numerical coordinates
@@ 5.2 $v_{B}$ from the entanglement wedge method
The numerical results of the butterfly velocity as a function of temperature and chemical potential are shown in Figure 8. To obtain this result, we numerically constructed approximately $\sim 5\times 10^{5}$ charged black hole solutions across a two-dimensional grid of initial conditions $\{\phi_{0},\Phi_{1}\}$. The parameter $\phi_{0}$ was varied in uniform increments from $0.55$ to $5.5$, while for each value of $\phi_{0}$, the ratio $\Phi_{1}/\Phi_{1}^{\text{max}}(\phi_{0})$ was varied from 0 to $0.5$ in equally spaced steps. [^5] For this system, the maximum value $\Phi_{1}$ can take is bounded by $\phi_{0}$ and is given by $\Phi_{1}<\Phi_{1}^{max}\sqrt{-\frac{2V(\phi_{0})}{f(\phi_{0})}}$. We again remind the reader that $\phi_{0}$ corresponds to the value of the scalar field at the horizon and $\Phi_{1}$ is the electric field at the horizon. We find that for relatively large temperatures, $T>200~\text{MeV}$, $v_{B}$ monotonically increases with the temperature, and then, as in the case of the other three models, it asymptotes to the conformal chargeless plasma value in the infinite-temperature limit. However, as opposed to the other three models, $v_{B}$ now also shows a nonmonotonic variation at relatively low temperature $T<130~\text{MeV}$, i.e., $v_{B}$ decreases with temperature for $T<130~\text{MeV}$.
At this point, it is important to keep in mind that, unlike the model considered in Section 4, there is no Hawking/Page phase transition in the gravity side in the current holographic model. This hinders the determination of the dual deconfinement transition temperature via direct methods in this model. To get the deconfinement temperature, one therefore relies on indirect methods, such as the temperature at which the speed of sound attains a minimum value. From this indirect method, the deconfinement transition temperature can be approximated to be around $T_{c}\sim 140~\text{MeV}$ [56]. Therefore, it is safe to say that the butterfly velocity again exhibits monotonic thermal structure in the high-temperature deconfined phase.
Furthermore, interestingly, just like in the case of the other three models, $v_{B}$ monotonically decreases with the chemical potential for all fixed values of the temperature. This observation further suggests a universal behavior in the rate of information propagation, associated with the spread of perturbations and operator growth, within the local degrees of freedom of the dual deconfined plasma, which decreases with increasing chemical potential. Note that, due to the marginal decrease in $v_{B}$ with respect to $\mu$, this fact is not clearly evident from Figure 8, but one can verify this by obtaining a two-dimensional plot of $v_{B}$ vs $\mu$ at various fixed values of $T$.
@@
However, the behavior at low temperatures, particularly near the deconfinement temperature, is less transparent. Unlike in the UV, there is no general symmetry argument that dictates the observed similarity between the butterfly velocity and the speed of sound in the infrared (IR). We also analyze other numerical holographic QCD models of the type [56], such as [55] and [106], and find analogous thermal behavior between $v_{B}$ and $c_{s}$. This apparent parallel thermal behavior may suggest deeper dynamical connections or shared sensitivity to the underlying nonperturbative physics of the strongly coupled plasma, though further investigation is required to elucidate the mechanisms responsible for this parallel behavior.
@@ 5.3 $v_{B}$ from the OTOC method
[…]
Computing the near-horizon limits of $A$ and $\partial_{\tilde{u}}\partial_{\tilde{v}}B$ as $\tilde{u}\rightarrow 0$, we arrive at the following:
@@ 5.4 $v_{B}$ from the pole-skipping method
Although analytic results for the energy-momentum tensor are difficult to obtain in the holographic QCD model considered in this section, it is straightforward to see that, as this model is diffeomorphism invariant and falls in the EMD gravity definition of type (3.1), the energy-momentum tensor again satisfies the condition (2.44). Accordingly, the pole-skipping points are again given by Eq. (2.47). Explicitly, after a little bit of algebra, they take the following form:
@@
$v_{B}^{2}$ $=$ $\frac{\omega_{*}^{2}}{q_{*}^{2}}=\frac{1}{6\eta^{\prime}(z_{h})\mathcal{G}_{0}^{\text{far}}}\,$ (5.24)
@@
We conclude by noting a technical subtlety in the holographic model considered in [56]. The last term in the gauge-kinetic function $\tilde{f}(\phi)$, as defined in Eq. (5.1), becomes effectively negligible for most values of $\phi$ due to the large numerical value of the parameter $b_{3}$, contributing meaningfully only in the limit of very small $\phi$. This exponential term is included to ensure that $\tilde{f}(0)=1$. However, this choice also influences the scaling of the chemical potential and charge, particularly when fitting to susceptibility data [from which the parameters of $\tilde{f}(\phi)$ are fixed], and should therefore be handled with care. Enforcing $\tilde{f}(0)=1$ results in a spurious narrow spike in the function $\tilde{f}(\phi)$ at small $\phi$. This issue has been discussed in great detail in [109], where it was further pointed out that the physical impact of this spike is essentially negligible.